# The Obstacle Problem Arising from the American Chooser Option

### v2

@@ 1. Introduction
(1.1) $\begin{cases}\partial_{t}V^{\rm ch}(t,s)+\mathfrak{L}V^{\rm ch}(t,s)\leqslant 0\quad\text{for }\ (t,s)\in D_{T}\ \text{ with }\ V^{\rm ch}(t,s)=\max\{C^{A}(t,s),P^{A}(t,s)\},\\ \partial_{t}V^{\rm ch}(t,s)+\mathfrak{L}V^{\rm ch}(t,s)=0\quad\text{for }\ (t,s)\in D_{T}\ \text{ with }\ V^{\rm ch}(t,s)>\max\{C^{A}(t,s),P^{A}(t,s)\},\\ V^{\rm ch}(T,s)=\max\{C^{A}(T,s),P^{A}(T,s)\}\quad\text{for }\ 0<s<\infty.\end{cases}$
@@
(1.2) $\begin{cases}\partial_{t}C^{A}(t,s)+\mathfrak{L}C^{A}(t,s)\leqslant 0\quad\text{for }\ (t,s)\in D_{T_{c}}\ \text{ with }\ C^{A}(t,s)=(s-K_{c})^{+},\\ \partial_{t}C^{A}(t,s)+\mathfrak{L}C^{A}(t,s)=0\quad\text{for }\ (t,s)\in D_{T_{c}}\ \text{ with }\ C^{A}(t,s)>(s-K_{c})^{+},\\ C^{A}(T,s)=(s-K_{c})^{+}\quad\text{for }\ 0<s<\infty.\end{cases}$
(1.3) $\begin{cases}\partial_{t}P^{A}(t,s)+\mathfrak{L}P^{A}(t,s)\leqslant 0\quad\text{for }\ (t,s)\in D_{T_{p}}\ \text{ with }\ P^{A}(t,s)=(K_{p}-s)^{+},\\ \partial_{t}P^{A}(t,s)+\mathfrak{L}P^{A}(t,s)=0\quad\text{for }\ (t,s)\in D_{T_{p}}\ \text{ with }\ P^{A}(t,s)>(K_{p}-s)^{+},\\ P^{A}(T,s)=(K_{p}-s)^{+}\quad\text{for }\ 0<s<\infty.\end{cases}$
@@
Standard American options have been extensively studied in the literature (e.g., [24, 26, 28, 15]). Most of these studies however rely on obstacle problems where the obstacle function has an explicit, closed-form representation. In contrast, the American chooser option involves an obstacle function that depends on the interaction between two separate obstacle problems. This interaction gives rise to a new level of mathematical difficulty, as the obstacle is no longer explicitly known but rather emerges as the maximum of two distinct solutions.
@@
A crucial aspect of our analysis is the structure of the free boundaries. Due to the nature of the American chooser option, the associated obstacle problem features two time-dependent free boundaries. Mathematically, we observe that the obstacle’s components align with solutions in separate, disjoint regions, as in [15]. However, unlike [15], the exact values of the obstacle at each point in the domain are unknown. The key insight is that these regions lie within the respective exercise regions of the American put and call options. This guarantees the existence of two distinct free boundaries provided that the exercise region of the American chooser option is nonempty. We further establish their monotonicity and smoothness, which are crucial for understanding the solution’s behavior.
Our paper is structured as follows. Section 2 provides the financial background and formal definition of the American chooser option and demonstrates the properties of the Amercian call and put options. Section 3 establishes the existence and uniqueness of the obstacle problem solution using the penalty method. Section 4 examines the monotonicity and smoothness of the two free boundaries that arise in this framework.
@@ 2.1. Notations
Throughout the paper, for each $\epsilon>0$, $\varphi_{\epsilon}$ is a smooth function in $\mathbb{R}$ satisfying that
(2.1) $\begin{cases}\varphi_{\varepsilon}\geqslant 0,\quad 0\leqslant\varphi^{\prime}_{\varepsilon}\leqslant 1,\quad\varphi^{\prime\prime}_{\varepsilon}\geqslant 0,\\ \varphi_{\varepsilon}(t)=t\ \ \text{if}\ \ t\geqslant\varepsilon,\quad\varphi_{\varepsilon}(t)=0\ \ \text{if}\ \ t\leqslant-\varepsilon,\\ \varphi_{\varepsilon}(t)\leqslant(t+\varepsilon)^{+}\ \ \text{for}\ \ t\in\mathbb{R},\\ \displaystyle{\lim_{\varepsilon\to 0}\varphi_{\varepsilon}(t)}=t^{+}\ \ \text{uniformly for }\ t\in\mathbb{R}.\end{cases}$
Let $\Omega_{T}^{n}:=(0,T)\times(-n,n)$.
  • $C^{k+\frac{\alpha}{2},2k+\alpha}(\Omega_{T}^{n})$, $\alpha\in(0,1)$, $k\in\mathbb{N}$ is the Banach space under the following norm for $V$:
    $\begin{split}&\|V\|_{C^{k+\frac{\alpha}{2},2k+\alpha}(\Omega_{T}^{n})}\\ &\quad:=\sum_{i+2j=0,2,\cdots,2k}\limits\sup_{(t_{1},x_{1}),(t_{2},x_{2})\in\Omega_{T}^{n}}\{|D^{i}_{x}D^{j}_{t}V(t_{1},x_{1})|+\frac{|D^{i}_{x}D^{j}_{t}V(t_{1},x_{1})-D^{i}_{x}D^{j}_{t}V(t_{2},x_{2})|}{{|t_{1}-t_{2}|}^{\frac{\alpha}{2}}+{|x_{1}-x_{2}|}^{\alpha}}\}.\end{split}$
  • $L^{p}(\Omega_{T}^{n}),p\geqslant 1$ is the completion of $C^{\infty}(\Omega_{T}^{n})$ under the following norm for $V$:
[…]
  • $W^{1,2}_{p}(\Omega_{T}^{n}),$ $p\geqslant 1$ is the completion of $C^{\infty}(\Omega_{T}^{n})$ under the following norm for $V$:
[…]
Let ${G}$ be a parabolic domain, $(\tau_{0},x_{0})\in G$ and $\varepsilon>0$.
  • $Q((\tau_{0},x_{0}),\varepsilon)$ is the cylinder such that
    $Q((\tau_{0},x_{0}),\varepsilon):=\{(\tau,x)\in\Omega_{T}:\text{max}\{|x-x_{0}|,|\tau-\tau_{0}|^{\frac{1}{2}}\}<\varepsilon,\tau<\tau_{0}\}.$
  • $\mathcal{P}{G}$ is the parabolic boundary of ${G}$ which is defined to be the set of all points $(\tau_{0},x_{0})\in\partial{G}$ such that for any $\varepsilon>0,$ the cylinder $Q((\tau_{0},x_{0}),\varepsilon)$ contains points not in ${G}$.
@@ 2.2. Financial Background: American Chooser Option
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a filtered probability space with a filtration $\{\mathcal{F}_{t}\}_{t\geqslant 0}$ satisfying the usual conditions. Under the risk-neutral measure $\mathbb{Q}$, the stock price $S_{t}$ follows a geometric Brownian motion (GBM) given by
@@
To define the American chooser option, we first introduce the American call and put options. These two options serve as the fundamental building blocks for the American chooser option. An American call option with strike price $K_{c}$ and expiration time $T_{c}$ gives the holder the right to buy the underlying asset at price $K_{c}$ at any time $\tau\in\mathcal{U}_{t,T_{c}}$, where $\mathcal{U}_{t,T}$ is the set of all $\mathcal{F}$-stopping times taking values in $[t,T]$. The value $C^{A}(t,S_{t})$ of the American call option at time $t$ is given by
(2.2) $C^{A}(t,S_{t})=\sup_{\tau\in\mathcal{U}_{t,T_{c}}}\mathbb{E}^{\mathbb{Q}}\left[e^{-r(\tau-t)}(S_{\tau}-K_{c})^{+}\mid\mathcal{F}_{t}\right].$
@@
(2.3) $P^{A}(t,S_{t})=\sup_{\tau\in\mathcal{U}_{t,T_{p}}}\mathbb{E}^{\mathbb{Q}}\left[e^{-r(\tau-t)}(K_{p}-S_{\tau})^{+}\mid\mathcal{F}_{t}\right].$
If $C^{A}$ and $P^{A}$ are strong solutions to the obstacle problems (1.2) and (1.3), respectively, then one can show, by applying Itô’s lemma for Sobolev space (see [17]), that they solve the corresponding optimal stopping problems (2.2) and (2.3); see, for instance, [3, Appendix B].
@@
(2.4) $V^{\rm ch}(t,S_{t})=\sup_{\tau\in\mathcal{U}_{t,T}}\mathbb{E}\left[e^{-r(\tau-t)}\max\left\{C^{A}(\tau,S_{\tau}),P^{A}(\tau,S_{\tau})\right\}\mid\mathcal{F}_{t}\right].$
Using dynamic programming and Itô’s lemma, we can easily derive the obstacle problem (1.1) from (2.4).
@@ 2.3. Change of variables
Since the obstacle problems in (1.2) and (1.3) are backward degenerate problems, we transform them into forward non-degenerate problems and denote the solutions by $\widehat{C}^{A}$ and $\widehat{P}^{A}$, respectively. More precisely, we introduce the following transformation:
@@
$\Omega_{\eta}:=\{(\zeta,x)\mid 0<\zeta<\eta,\;x\in\mathbb{R}\}$
@@
According to the vast literature on American options (see, for instance, [1, 15, 16, 20, 25, 27]), the obstacle problems (2.5) and (2.6) have unique strong solutions, $\widehat{C}^{A}\in W^{1,2}_{p\textup{,}\,\text{loc}}(\Omega_{T_{c}})\cap C(\overline{\Omega_{T_{c}}})$ and $\widehat{P}^{A}\in W^{1,2}_{p\textup{,}\,\text{loc}}(\Omega_{T_{p}})\cap C(\overline{\Omega_{T_{p}}})$, respectively.
@@
Then, $V(\tau,x)$ satisfies the following forward non-degenerate obstacle problems:
@@
(2.8) $\begin{cases}C(\tau,x):=\widehat{C}^{A}(T_{c}-T+\tau,x)=C^{A}(T-\tau,e^{x}),\\ P(\tau,x):=\widehat{P}^{A}(T_{p}-T+\tau,x)=P^{A}(T-\tau,e^{x}),\end{cases}\quad\tau\in[0,T].$
@@ 2.4. Properties of American Call and Put Options
In each obstacle problem in (2.5) and (2.6), we define the exercise regions, denoted by $\mathcal{E}_{C}$ and $\mathcal{E}_{P}$, and the continuation regions, denoted by $\mathcal{C}_{C}$ and $\mathcal{C}_{P}$, as follows:
@@
It is well known that $\hat{x}_{c}(\tau)$ and $\hat{x}_{p}(\tau)$ satisfy the following properties (see [2, 4, 8, 21, 27]):
@@
For further discussion, it is necessary to verify various properties of ${C}$ and ${P}$. Since ${C}$ and ${P}$ are restricted functions of $\widehat{C}^{A}$ and $\widehat{P}^{A}$, it is natural to analyze $\widehat{C}^{A}$ and $\widehat{P}^{A}$ over their original whole domains and then conclude that the same properties hold for ${C}$ and ${P}$, respectively. The following results are essentially well-known and their proofs are provided in the Appendix.
@@
Let $\widehat{C}^{A}$ and $\widehat{P}^{A}$ be the solution to (2.5) and (2.6), respectively.
  • (i)
    $20\leqslant\widehat{C}^{A}(\zeta,x)\leqslant e^{x}+$ 2 in $\Omega_{T_{c}}$, and $20\leqslant\widehat{P}^{A}(\zeta,x)\leqslant K_{p}+$ 2 in $\Omega_{T_{p}}$. Hence, by (2.8), we have that for all $(\tau,x)\in\Omega_{T}$,
    $20\leqslant C(\tau,x)\leqslant e^{x}+2\quad\text{and}\quad 0\leqslant P(\tau,x)\leqslant K_{p}+$ 2 .
  • (ii)
    $0\leqslant\partial_{x}\widehat{C}^{A}(\zeta,x)\leqslant e^{x}$ in $\Omega_{T_{c}}$, and $-e^{x}\leqslant\partial_{x}\widehat{P}^{A}(\zeta,x)\leqslant 0$ in $\Omega_{T_{p}}$. Hence, by (2.8), we have that for all $(\tau,x)\in\Omega_{T}$,
    $0\leqslant\partial_{x}C(\tau,x)\leqslant e^{x}\quad\text{and}\quad-e^{x}\leqslant\partial_{x}P(\tau,x)\leqslant 0$
  • (iii)
    $\partial_{\tau}\widehat{C}^{A}\geqslant 0$ in $\Omega_{T_{c}}$ and $\partial_{\tau}\widehat{P}^{A}\geqslant 0$ in $\Omega_{T_{p}}$. Hence, by (2.8), we have that for all $(\tau,x)\in\Omega_{T}$,
    $\partial_{\tau}C(\tau,x)\geqslant 0\quad\text{and}\quad\partial_{\tau}P(\tau,x)\geqslant 0.$
  • (iv)
    For each $\tau\in(0,T]$, there exists $x_{\tau}\in\mathbb{R}$ such that $C(\tau,x)>P(\tau,x)$ for all $x\geqslant x_{\tau}$.
Proof.
See the Appendix. ∎

### v3

@@ 1. Introduction
[…]
@@
$D_{\eta}:=\{(t,s)\;:\;0<t<\eta,\ 0<s<\infty\}$ for $\eta>0$, and the functions $C^{A}$ and $P^{A}$ satisfy the following obstacle problems, respectively:
@@
[…]
@@
Standard American options have been extensively studied in the literature (e.g., [28, 29, 27, 12]). Most of these studies however rely on obstacle problems where the obstacle function has an explicit, closed-form representation. In contrast, the American chooser option involves an obstacle function that depends on the interaction between two separate obstacle problems. This interaction gives rise to a new level of mathematical difficulty, as the obstacle is no longer explicitly known but rather emerges as the maximum of two distinct solutions. More precisely, unlike the existing literature, our obstacle is given by the maximum of two strong solutions to separate obstacle problems. Therefore, one cannot directly apply the operator to the obstacle and compute it in a pointwise manner as in standard problems with an explicit payoff, and this feature makes the analysis highly challenging.
@@
A crucial part of our analysis concerns the structure of the free boundaries. Owing to the nature of the American chooser option, the associated obstacle problem gives rise to two time-dependent free boundaries. More precisely, the two components of the obstacle coincide with the relevant solutions on two separate and disjoint regions, similarly to [12]. However, unlike [12], the obstacle is not explicitly known pointwise throughout the domain. The key observation is that these regions are contained in the respective exercise regions of the American put and call options. As a consequence, contact occurs only along the smooth payoff parts $e^{x}-K_{c}$ and $K_{p}-e^{x}$ of the obstacle, which allows us to employ standard arguments to establish the monotonicity and smoothness of the free boundaries.
Our paper is structured as follows. Section 2 provides the financial background and formal definition of the American chooser option and demonstrates the properties of the American call and put options. Section 3 establishes the existence and uniqueness of the obstacle problem solution using the penalty method. Section 4 examines the monotonicity and smoothness of the two free boundaries that arise in this framework.
@@ 2.1. Notations
Throughout the paper, for each $\varepsilon>0$, $\varphi_{\varepsilon}\in C^{\infty}(\mathbb{R})$ is a function that satisfies
(2.1) $\begin{cases}\varphi_{\varepsilon}\geqslant 0,\quad 0\leqslant\varphi^{\prime}_{\varepsilon}\leqslant 1,\quad\varphi^{\prime\prime}_{\varepsilon}\geqslant 0,\\ \varphi_{\varepsilon}(\lambda)=\lambda\ \ \text{if}\ \ \lambda\geqslant\varepsilon,\quad\varphi_{\varepsilon}(\lambda)=0\ \ \text{if}\ \ \lambda\leqslant-\varepsilon,\\ \varphi_{\varepsilon}(\lambda)\leqslant(\lambda+\varepsilon)^{+}\ \ \text{for}\ \ \lambda\in\mathbb{R},\\ \displaystyle{\lim_{\varepsilon\to 0}\varphi_{\varepsilon}(\lambda)}=\lambda^{+}\ \ \text{uniformly for }\ \lambda\in\mathbb{R}.\end{cases}$
Let $\mathcal{D}\subset\mathbb{R}^{d+1}$ be a bounded parabolic domain, where $d\in\mathbb{N}$ denotes the dimension of the spatial variable $x$ and $t$ represents the time variable.
  • $C^{k+\frac{\alpha}{2},2k+\alpha}(\mathcal{D})$, $\alpha\in(0,1)$, $k\in\mathbb{N}$ is the Banach space under the following norm for $V$:
    $\begin{split}&\|V\|_{C^{k+\frac{\alpha}{2},2k+\alpha}(\mathcal{D})}\\ &\quad:=\sum_{i+2j=0,2,\cdots,2k}\limits\sup_{(t_{1},x_{1}),(t_{2},x_{2})\in\mathcal{D}}\{|D^{i}_{x}D^{j}_{t}V(t_{1},x_{1})|+\frac{|D^{i}_{x}D^{j}_{t}V(t_{1},x_{1})-D^{i}_{x}D^{j}_{t}V(t_{2},x_{2})|}{{|t_{1}-t_{2}|}^{\frac{\alpha}{2}}+{|x_{1}-x_{2}|}^{\alpha}}\}.\end{split}$
  • $L^{p}(\mathcal{D}),p\geqslant 1$ is the completion of $C^{\infty}(\mathcal{D})$ under the following norm for $V$:
    $\|V\|_{L^{p}(\mathcal{D})}:=\left(\iint_{\mathcal{D}}{|V(t,x)|}^{p}\,dx\,dt\right)^{\frac{1}{p}}.$
  • $W^{1,2}_{p}(\mathcal{D}),$ $p\geqslant 1$ is the completion of $C^{\infty}(\mathcal{D})$ under the following norm for $V$:
    $\|V\|_{W^{1,2}_{p}(\mathcal{D})}:=\left(\iint_{\mathcal{D}}\big\{|V|^{p}+|\partial_{t}V|^{p}+|\partial_{x}V|^{p}+|\partial^{2}_{x}V|^{p}\big\}\,dx\,dt\right)^{\frac{1}{p}}.$
Let $(t_{0},x_{0})\in\mathcal{D}$ and $r>0$.
  • $Q((t_{0},x_{0}),r)$ is the cylinder such that
    $Q((t_{0},x_{0}),r):=\{(t,x)\in\mathcal{D}:\text{max}\{|x-x_{0}|,|t-t_{0}|^{\frac{1}{2}}\}<r,t<t_{0}\}.$
  • $\mathcal{P}\mathcal{D}$ is the parabolic boundary of $\mathcal{D}$ which is defined to be the set of all points $(t_{0},x_{0})\in\partial{\mathcal{D}}$ such that for any $r>0,$ the cylinder $Q((t_{0},x_{0}),r)$ contains points not in $\mathcal{D}$}.
  • $\mathcal{BD}$ is the set of all points $(t,x)\in\mathcal{P}\mathcal{D}$ such that there is a positive $r>0$ with
    $Q((t+r^{2},x),r)\subset\mathcal{D}.$
@@ 2.2. Financial Background: American Chooser Option
Let $(\Omega,\mathcal{G},\mathbb{P})$ be a filtered probability space with a filtration $\{\mathcal{G}_{t}\}_{t\geqslant 0}$ satisfying the usual conditions. Under the risk-neutral measure $\mathbb{Q}$, the stock price $S_{t}$ follows a geometric Brownian motion (GBM) given by
@@
To define the American chooser option, we first introduce the American call and put options. These two options serve as the fundamental building blocks for the American chooser option. An American call option with strike price $K_{c}$ and expiration time $T_{c}$ gives the holder the right to buy the underlying asset at price $K_{c}$ at any time $\theta\in\mathcal{U}_{t,T_{c}}$, where $\mathcal{U}_{t,T}$ is the set of all $\mathcal{G}$-stopping times taking values in $[t,T]$. The value $C^{A}(t,S_{t})$ of the American call option at time $t$ is given by
(2.2) $C^{A}(t,S_{t})=\sup_{\theta\in\mathcal{U}_{t,T_{c}}}\mathbb{E}^{\mathbb{Q}}\left[e^{-r(\theta-t)}(S_{\theta}-K_{c})^{+}\mid\mathcal{G}_{t}\right].$
@@
(2.3) $P^{A}(t,S_{t})=\sup_{\theta\in\mathcal{U}_{t,T_{p}}}\mathbb{E}^{\mathbb{Q}}\left[e^{-r(\theta-t)}(K_{p}-S_{\theta})^{+}\mid\mathcal{G}_{t}\right].$
If $C^{A}$ and $P^{A}$ are strong solutions to the obstacle problems (1.2) and (1.3), respectively, then one can show, by applying Itô’s lemma for Sobolev spaces (see [14]), that they solve the corresponding optimal stopping problems (2.2) and (2.3); see, for instance, [5, Appendix B].
@@
(2.4) $V^{\rm ch}(t,S_{t})=\sup_{\theta\in\mathcal{U}_{t,T}}\mathbb{E}\left[e^{-r(\theta-t)}\max\left\{C^{A}(\theta,S_{\theta}),P^{A}(\theta,S_{\theta})\right\}\mid\mathcal{G}_{t}\right].$
Using dynamic programming and Itô’s lemma, we can easily derive the obstacle problem (1.1) from (2.4) (see Appendix A.).
@@ 2.3. Change of variables
We transform them into forward non-degenerate problems and denote the solutions by $\widehat{C}^{A}$ and $\widehat{P}^{A}$, respectively. More precisely, we introduce the following transformation:
@@
$\Omega_{\eta}:=\{(\zeta,x)\mid 0<\zeta<\eta,\;x\in\mathbb{R}\}.$
@@
According to the vast literature on American options (see, for instance, [1, 12, 13, 19, 24, 26, 27]), the obstacle problems (2.5) and (2.6) have unique strong solutions, $\widehat{C}^{A}\in W^{1,2}_{p\textup{,}\,\text{loc}}(\Omega_{T_{c}})\cap C(\overline{\Omega_{T_{c}}})$ and $\widehat{P}^{A}\in W^{1,2}_{p\textup{,}\,\text{loc}}(\Omega_{T_{p}})\cap C(\overline{\Omega_{T_{p}}})$, respectively.
@@
Then, $V(\tau,x)$ satisfies the following forward non-degenerate obstacle problem:
@@
(2.8) $\begin{cases}C(\tau,x):=\widehat{C}^{A}(T_{c}-T+\tau,x)=C^{A}(T-\tau,e^{x}),\\ P(\tau,x):=\widehat{P}^{A}(T_{p}-T+\tau,x)=P^{A}(T-\tau,e^{x}),\end{cases}\quad(\tau,x)\in\overline{\Omega_{T}}.$
@@ 2.4. Properties of American Call and Put Options
In each obstacle problem (2.5) and (2.6), it is well known that contact with the respective obstacles occurs only at their positive parts. Thus, we define the exercise regions, denoted by $\mathcal{E}_{C}$ and $\mathcal{E}_{P}$, and the continuation regions, denoted by $\mathcal{C}_{C}$ and $\mathcal{C}_{P}$, as follows:
@@
It is well known that $\hat{x}_{c}(\tau)$ and $\hat{x}_{p}(\tau)$ satisfy the following properties (see [2, 4, 6, 20, 26, 27]):
@@
For further discussion, it is necessary to verify various properties of C{C} and P{P}. Since C{C} and P{P} are restricted functions of $\widehat{C}^{A}$ and $\widehat{P}^{A}$, it is natural to analyze $\widehat{C}^{A}$ and $\widehat{P}^{A}$ over their original whole domains and then conclude that the same properties hold for C{C} and P{P}, respectively. The following results are essentially well-known and their proofs are provided in the Appendix.
@@
Let $\widehat{C}^{A}$ and $\widehat{P}^{A}$ be the solution to (2.5) and (2.6), respectively, and define for each $\lambda\in\mathbb{R}$ that $\Omega_{\lambda}:=(0,\lambda)\times\mathbb{R}$. Then the following properties hold.
  1. (i)
    $0\leqslant\widehat{C}^{A}\leqslant e^{x}+2$ in $\Omega_{T_{c}}$,  $0\leqslant\widehat{P}^{A}\leqslant K_{p}+2$ in $\Omega_{T_{p}}$.
  2. (ii)
    $0\leqslant\partial_{x}\widehat{C}^{A}\leqslant e^{x}$ in $\Omega_{T_{c}}$,  $e^{x}\leqslant\partial_{x}\widehat{P}^{A}\leqslant 0$ in $\Omega_{T_{p}}$.
  3. (iii)
    $\partial_{\zeta}\widehat{C}^{A}\geqslant 0$ in $\Omega_{T_{c}}$,  $\partial_{\zeta}\widehat{P}^{A}\geqslant 0$ in $\Omega_{T_{p}}$.
  4. (iv)
    For each $\tau\in(0,T]$, there exists $x_{\tau}\in\mathbb{R}$ such that $C(\tau,x)>P(\tau,x)$ for all $x>x_{\tau}$.
Moreover, by (2.8), the corresponding statements hold for $C$ and $P$ in $\Omega_{T}$.